OpenStudy (anonymous):

If we have the following sequence, find the n-th term 2,3,5,7,11,13,...

OpenStudy (dmezzullo):

this is the sequence of prime numbers. you can use lagrange interpolation though, if not necessarily prime numbers

OpenStudy (anonymous):

yup, can I say that the n-th prime is bounded?

Parth (parthkohli):

There is no closed form for primes.

OpenStudy (anonymous):

where a,b are arbitrary constants

OpenStudy (anonymous):

\[anlogn \le p _{n} \le bnlogn\]

Parth (parthkohli):

no, this is not a sequence generator.

OpenStudy (anonymous):

where \[n \in N, n \ge 2\]